figure out the y-intercept of the graph of the equation y=x2 (this is x squared) 6x = 9 and -3/2 (x-6)2 (squared).
Answer by stanbon(75887)

According to Newton's Law of Gravitation the force of attraction between two objects varies jointly as their masses and inversely square of the distance between them. How does the force change if each of their masses is tripled but the distance between them is cut to one-third?
---------------
a = k(m1*m2/d^2)
----------
After change:
a = k(3m1*3m2/((1/3)d)^2
a = k(9m1*m2/((1/9)d^2)
a = 81k(m1*m2/d^2)
===============
Cheers,
stan H.